(1). What is the final velocity of a bike that starts at 4 m/s and has zero  acceleration act on it for 10 seconds?

(2).How far will the bike in question 1 travel in the 10-second time period?



1
Expert's answer
2021-01-01T14:54:39-0500

1)


"v_f=v_i+at\\v_f=4+0(10)=4\frac{m}{s}"

2)


"s=vt=4(10)=40\ m"